# ManiCM: Real-time 3D Diffusion Policy via Consistency Model for Robotic Manipulation

### v1

@@ Abstract
Diffusion models have been verified to be effective in generating complex distributions from natural images to motion trajectories. Recent diffusion-based methods show impressive performance in 3D robotic manipulation tasks, whereas they suffer from severe runtime inefficiency due to multiple denoising steps, especially with high-dimensional observations. To this end, we propose a real-time robotic manipulation model named ManiCM that imposes the consistency constraint to the diffusion process, so that the model can generate robot actions in only one-step inference. Specifically, we formulate a consistent diffusion process in the robot action space conditioned on the point cloud input, where the original action is required to be directly denoised from any point along the ODE trajectory. To model this process, we design a consistency distillation technique to predict the action sample directly instead of predicting the noise within the vision community for fast convergence in the low-dimensional action manifold. We evaluate ManiCM on 31 robotic manipulation tasks from Adroit and Metaworld, and the results demonstrate that our approach accelerates the state-of-the-art method by 10 times in average inference speed while maintaining competitive average success rate.
1 Introduction
@@
Designing robots for diverse manipulation tasks has been highly desired in the robotic community for a long time. In this pursuit, previous arts have made great progress in exploring different architectures, e.g., perceptive models like convolutional networks [34, 15], transformers [4, 3], and generative models like diffusion models [7, 48]. Among these choices, diffusion-based policies receive increasing attention for their capacity to model diverse high-dimensional robotic trajectories, where robot observation is provided as the condition for generation. For instance, diffusion models have been extensively utilized across various domains as high-level procedure planners [16, 1, 11, 26, 12], low-level motion policies [29, 41, 7, 42, 24, 44], teacher models [21, 47], and data synthesis engines [20, 46, 13], demonstrating impressive performance in diverse robotic tasks.
@@
Figure 1: Trade-off Between Efficiency and Effectiveness. We present ManiCM, a real-time 3D diffusion policy by imposing the consistency constraint to the diffusion process. BCRNN+3D [25] is a state-of-the-art perceptive model-based behavior cloning agent augmented with 3D point cloud input. DP [7] and DP3 [48] are the state-of-the-art diffusion-based manipulation agents. ManiCM achieves a decision-making runtime of 17ms, which is 10$\times$ faster than previous mainstream methods.
@@
However, diffusion-based fashions inevitably suffer from severe runtime inefficiency by requiring considerable sampling steps for high-quality action synthesis during inference, even with some sampling acceleration [36]. As shown in Figure 1, the diffusion policy (DP [7]) requires $\sim$162ms per step to generate a high-quality action. Moreover, the 3D diffusion policy (DP3 [48]) only achieves a decision latency of $\sim$178ms when dealing with the high-dimensional 3D point cloud input, which blocks the applications of real-time closed-loop control. To address this issue, recent efforts are proposed to speed up the inference phase of diffusion models by hierarchical sampling [10, 42, 24]. Nevertheless, it is hard to determine the hierarchy across task domains that exhibit different challenging levels, which limits the practicality of the existing works to 3D robotic manipulation.
@@
Recently, the concept of consistency models [37, 22] has been introduced in image generation, which synthetize high-fidelity images with a minimal number of steps. In this paper, we propose a ManiCM method that leverages the concept of consistency models for real-time robotic manipulation. Different from existing methods that focus on designing hierarchical planning for specific domains, we design a generalizable consistency distillation technique to generate robot actions in only one-step inference. More specifically, we first formulate a diffusion process to denoise robot action conditioned on point cloud, and then ensure the consistency property that the original action can be directly denoised from any point along the ODE trajectory. Then, we utilize the consistency distillation technique to predict the action sample directly, based on the observation that predicting the denoised action converges faster than predicting the noise in the vision community. We evaluate ManiCM on a diverse set of robotic manipulation tasks from Adroit and Metaworld, and the results illustrate that our approach achieves an average inference speed acceleration of 10$\times$ compared to the state-of-the-art method, while still maintaining competitive average success rates (Figure 1). We introduce consistency distillation into the robotic manipulation area and accelerate high-quality 3D manipulation action generation to a real-time level. Our contributions can be concluded as follows:
@@
  • We propose a real-time 3D diffusion policy to learn robot action conditioned on point cloud, so that the original action can be directly denoised from any point along the ODE trajectory.
@@
2 Related Work
2.1 Diffusion Models for 3D Robotic Manipulation
@@
The exploration of the most effective architecture for end-to-end robotic manipulation has been a long-standing process. The trended architectures include convolution networks [34, 15], transformers [4, 3], and diffusion models [16, 1, 11, 26, 12, 29, 41, 7, 20, 46, 13, 48, 48, 24, 44]. Among these, diffusion models have first emerged as a transformative approach in image generation [14, 36], leveraging the iterative refinement of Gaussian noise into data samples. Recently, they have been introduced to the field of embodied AI, and have made significant progress in applications that demand nuanced decision-making and adaptive control. Specifically, [7, 16, 1] first demonstrate the potential of the diffusion model in low-dimensional control tasks. Furthermore, the diffusion model classes [48, 17, 24, 44] are extended to more complex 3D robotic manipulation tasks, and also achieve superior performance than other traditional architectures. However, applying diffusion models to 3D robotic manipulation faces the issue of insufficient decision-making frequency. 3D tasks involve complex spatial representations, which pose significant challenges for real-time robotic operation and control. Denoising processes are effective at generating high-fidelity outputs, but their iterative nature can be computationally expensive, creating a conflict between thorough exploration of the solution space and the need for fast decision-making in dynamic robotic tasks. To solve this issue, recent methods present to skip inference steps by hierarchical sampling [10, 42, 24], while it is hard to determine the hierarchy across different domains. In this paper, we harness the power of the consistency model [37] to accelerate the efficiency of diffusion models in 3D decision-making, utilizing point clouds as effective 3D representations.
2.2 Consistency Models
@@
To accelerate the sampling speed of the current diffusion models (DMs) [35, 14, 27, 32], the concept of consistency models (CMs) [37] is first introduced in the image generation domain. The core idea behind the consistency models is to enforce the model to learn to map any point at any time to the origin of the probability flow ODE (PF-ODE) trajectory, i.e., the clean image. The consistency model facilitates efficient image generation, allowing trade-offs between extremely few inference steps and the generation quality. Recently, latent consistency models (LCMs) [22, 23] successfully distilled from the Stable Diffusion [32], achieving significant acceleration in the speed of text-conditioned image generation. Based on LCMs, [5, 43] introduce diverse conditional controls to manipulate the text-to-image latent consistency models. Moreover, the consistent model has been introduced in various research domains such as video [40, 39], 3D human motion [8], and audio [2], unlocking new capabilities for various real-time applications. While initial success has been achieved in incorporating relatively simple embodied tasks with the consistency model, as explored in [6, 30, 9], its application to complex robotic manipulation tasks with high-dimensional 3D visual conditions still remain largely unexplored. To fill this research gap, we propose ManiCM to achieve real-time 3D diffusion policy generation for the first time.
@@
3 Approach
In this section, we first present the problem formulation and the preliminaries on consistency models (Section 3.1). Then, we present an overview of our pipeline, and detail the design of each component in the manipulation consistency model (Section 3.2). To enable our model to learn consistent generation from the teacher network, we propose the manipulation consistency distillation (Section 3.3).
3.1 Problem Formulation
The demand for robotic manipulation is of great significance in the pursuit of the general embodied agent. To train a manipulation agent, we focus on imitation learning that learns a policy from a limited set of expert demonstrations. The policy $\pi:\mathcal{O}\longmapsto\mathcal{A}$ maps the observation $o\in\mathcal{O}$ to the robot action $\boldsymbol{a}\in\mathcal{A}$. The observation contains the point cloud received from an eye-to-hand RGB-D camera and the robot proprioception data. The action space varies among different tasks and robot settings, which typically demonstrates the SE(3) motion of the end-effector followed by a normalized torque that the gripper fingers should apply. Existing methods leverage powerful diffusion models to learn the complex distribution in manipulation tasks, thereby generating high-quality robot actions. However, diffusion models demand considerable runtime during the inference phase due to the costly iterative sampling, leading to low decision frequency in real-time deployments especially with high-dimensional 3D visual inputs. Therefore, we propose a framework based on the consistency model, enabling to perform inference in only a single step, hence significantly enhancing decision efficiency for complex 3D manipulation tasks.
@@ 3.1 Problem Formulation
As shown in Equation 1, the implication of the self-consistency property is that for any input pairs $(\mathbf{x}_{t},t)$ on the same PF-ODE trajectory, their outputs $\textbf{{{f}}}(\mathbf{x}_{t},t)$ remain consistent. All consistency models have to meet the boundary condition $\textbf{{{f}}}(\mathbf{x}_{\epsilon},\epsilon)=\mathbf{x}_{\epsilon}$, as it plays an important role in the success of consistency model training, according to [37]. To impose the self-consistency property to deep neural networks, we implement function $\textbf{{f}}(\cdot,\cdot)$ as:
@@
where $d(\cdot,\cdot)$ is a metric function chosen for measuring the distance between two samples. $\textbf{{{f}}}_{\bf\theta}(\cdot,\cdot)$ and $\textbf{{{f}}}_{{\bf\theta}^{-}}(\cdot,\cdot)$ are referred to as ‘online network’ and ‘target network’ according to [37]. When the optimization of the online network converges, the target network will eventually match the online network since $\bf{\theta}^{-}$ is a running average of $\bf{\theta}$. The estimated consistency model can become arbitrarily accurate as long as the step size of ODE solver is sufficiently small and the consistency distillation loss reaches zero. Importantly, the boundary condition $\textit{f}_{\theta}(x,\epsilon)\equiv x$ ensures that the model does not converge to a meaningless solution $\textit{f}_{\theta}(x,t)\equiv 0$ in training.
3.2 ManiCM: Manipulation Consistency Model
The overall pipeline of our ManiCM is shown in Figure 2, in which we construct a manipulation consistency model to tackle the low decision efficiency in 3D manipulation diffusion models [48], releasing the potential of Consistency Models to enhance the manipulation generation task. We choose the powerful 3D Diffusion Policy (DP3 [48]) as the underlying diffusion model to distill its knowledge into a single-step sampler to achieve real-time inference without compromising manipulation quality. DP3 integrates 3D point cloud conditions into the Diffusion Policy (DP [7]), leveraging both the spatial understanding capabilities intrinsic to 3D modalities and the expressive power of the diffusion policy. In the training phase, ManiCM implements a manipulation consistency distillation to distill the knowledge from the teacher diffusion-based policy. We adopt a manipulation consistency function to predict the action sample, opposite to the noise prediction that is commonly employed in image generation. This modification results in a faster convergence to the low-dimensional robot action manifold. In the evaluation phase, ManiCM is able to decode high-quality action within one inference.
3D Conditional Generation. To incorporate the rich spatial information embedded in the 3D point cloud, we proposes to inject the point cloud representation into the diffusion policy as conditions. Specifically, given the depth image received from a single-view RGB-D camera, we first obtain the point cloud by the extrinsic and intrinsic of the camera. We down-sample the point cloud via farthest point sampling, in order to focus on the interest region like the robot arm and targets. Then, the sparse point cloud is encoded by a carefully designed MLP point cloud encoder, resulting in a compact 3D representation. Subsequently, we condition the original diffusion model with the compact 3D representation, which can be expressed as:
@@ 3.2 ManiCM: Manipulation Consistency Model
where $\boldsymbol{\epsilon}_{\theta}$ is the predicted noise. $p$ and $q$ denote point cloud conditions and robot poses, respectively. Equation 4 is also known as the reverse process [14]. By incorporating point cloud conditions, the model can capture the spatial information that helps fine-grained robotic manipulation tasks.
@@
where $\textbf{{{F}}}_{\mathbf{\theta}}$ is a manipulation consistency function that contains the accumulated $\boldsymbol{\epsilon}_{\theta}$ in Equation 4 along the trajectory. In practice, $c_{\text{skip}}(t)=\sigma_{d}^{2}/(\beta^{2}t^{2}+\sigma_{d}^{2})$, $c_{\text{skip}}(t)=\beta t/\sqrt{(\beta^{2}t^{2}+\sigma_{d}^{2})}$, where $\beta$ denotes timestep scaling value while $\sigma_{d}$ is a balance value. The $\textbf{{{F}}}_{\mathbf{\theta}}(\cdot,\cdot)$ can be parametrized to predict the action sample or the noise epsilon. In general cases, both prediction methods are viable because they are essentially optimizing the same diffusion process. In mainstream diffusion generation works for image generation, the default prediction type is noise rather than sample. We speculate that while directly predicting samples may intuitively yield higher-quality generated outputs [18], predicting the distribution of images is much more challenging for denoising models compared to predicting the distribution of noise in high-dimensional image space. However, the dimensionality of the robot action is quite small (approximately $28$) in manipulation tasks. In this scenario, directly learning the distribution of the samples is relatively easy and brings low variance and fast convergence, echoing [48]. More importantly, the large prediction variance caused by predicting noise will be amplified in consistency models. Specifically, the consistency property encourage the $\textbf{{{f}}}_{{\bf\theta}}(\boldsymbol{a}_{n},t_{n})$ to match $\hat{\boldsymbol{a}}_{0}=\textbf{{{f}}}_{{\bf\theta}}(\boldsymbol{a}_{0},t_{0})$, while the original diffusion model only requires $\hat{\boldsymbol{a}}_{n}$ to be close to $\hat{\boldsymbol{a}}_{n-1}$. Thus the variance is propagated along the ODE trajectory in consistency modeling, which ultimately blocks the convergence of prediction noise in manipulation tasks (Figure 3). Therefore, we parameterize the manipulation self-consistency function $\textbf{{{F}}}_{\mathbf{\theta}}(\cdot,\cdot)$ to predict the action sample rather than noise.
3.3 Manipulation Consistency Distillation

### v2

@@ Abstract
Diffusion models have been verified to be effective in generating complex distributions from natural images to motion trajectories. Recent diffusion-based methods show impressive performance in 3D robotic manipulation tasks, whereas they suffer from severe runtime inefficiency due to multiple denoising steps, especially with high-dimensional observations. To this end, we propose a real-time robotic manipulation model named ManiCM that imposes the consistency constraint on the diffusion process, so that the model can generate robot actions in only one-step inference. Specifically, we formulate a consistent diffusion process in the robot action space conditioned on the point cloud input, where the original action is required to be directly denoised from any point along the ODE trajectory. To model this process, we design a consistency distillation technique to predict the action sample directly instead of predicting the noise within the vision community for fast convergence in the low-dimensional action manifold. We evaluate ManiCM on 31 robotic manipulation tasks from Adroit and Metaworld, and the results demonstrate that our approach accelerates the state-of-the-art method by 10 times in the average inference speed while maintaining competitive average success rate.
I Introduction
@@
Designing robots for diverse manipulation tasks has been highly desired in the robotic community for a long time. In this pursuit, previous arts have made great progress in exploring different architectures, e.g., perceptive models like convolutional networks [1, 2], transformers [3, 4], and generative models like diffusion models [5, 6, 7]. Among these choices, diffusion-based policies receive increasing attention for their capacity to model diverse high-dimensional robotic trajectories, where robot observation is provided as the condition for generation. For instance, diffusion models have been extensively utilized across various domains as high-level procedure planners [8, 9, 10, 11, 12], low-level motion policies [13, 14, 5, 15, 16, 17], teacher models [18, 19], and data synthesis engines [20, 21, 22], demonstrating impressive performance in diverse robotic tasks.
@@
Figure 1: Trade-off Between Efficiency and Effectiveness. We present ManiCM, a real-time 3D diffusion policy by imposing the consistency constraint on the diffusion process. BCRNN+3D [23] is a state-of-the-art perceptive model-based behavior cloning agent augmented with 3D point cloud input. DP [5] and DP3 [6] are the state-of-the-art diffusion-based manipulation agents. ManiCM achieves a decision-making runtime of 17ms, which is 10$\times$ faster than previous mainstream methods.
@@
However, diffusion-based fashions inevitably suffer from severe runtime inefficiency by requiring considerable sampling steps for high-quality action synthesis during inference, even with some sampling acceleration [24]. As shown in Figure 1, the diffusion policy (DP [5]) requires $\sim$162ms per step to generate a high-quality action. Moreover, the 3D diffusion policy (DP3 [6]) only achieves a decision latency of $\sim$178ms when dealing with the high-dimensional 3D point cloud input, which blocks the applications of real-time closed-loop control. To address this issue, recent efforts have been proposed to speed up the inference phase of diffusion models by hierarchical sampling [25, 15, 16]. Nevertheless, it is hard to determine the hierarchy across task domains that exhibit different challenging levels, which limits the practicality of the existing works to 3D robotic manipulation.
@@
Recently, the concept of consistency models [26, 27] has been introduced in image generation, which synthetizes high-fidelity images with a minimal number of steps. In this paper, we propose a ManiCM method that leverages the concept of consistency models for real-time robotic manipulation. Different from existing methods that focus on designing hierarchical planning for specific domains, we design a generalizable consistency distillation technique to generate robot actions in only one-step inference. More specifically, we first formulate a diffusion process to denoise robot action conditioned on the point cloud, and then ensure the consistency property that the original action can be directly denoised from any point along the ODE trajectory. Then, we utilize the consistency distillation technique to predict the action sample directly, based on the observation that predicting the denoised action converges faster than predicting the noise in the vision community. We evaluate ManiCM on a diverse set of robotic manipulation tasks from Adroit and Metaworld, and the results illustrate that our approach achieves an average inference speed acceleration of 10$\times$ compared to the state-of-the-art method, while still maintaining competitive average success rates (Figure 1). We introduce consistency distillation into the robotic manipulation area and accelerate high-quality 3D manipulation action generation to a real-time level. Our contributions can be concluded as follows:
@@
  • We propose a real-time 3D diffusion policy to learn robot action conditioned on the point cloud, so that the original action can be directly denoised from any point along the ODE trajectory.
@@
II Related Work
II-A Diffusion Models for 3D Robotic Manipulation
@@
The exploration of the most effective architecture for end-to-end robotic manipulation has been a long-standing process. The trended architectures include convolution networks [1, 2], transformers [3, 4], and diffusion models [8, 9, 10, 11, 12, 13, 14, 5, 20, 21, 22, 6, 16, 17]. Among these, diffusion models first emerged as a transformative approach in image generation [28, 24], leveraging the iterative refinement of Gaussian noise into data samples. Recently, they have been introduced to the field of embodied AI, and have made significant progress in applications that demand nuanced decision-making and adaptive control. Specifically, [5, 8, 9] first demonstrated the potential of the diffusion model in low-dimensional control tasks. Furthermore, the diffusion model classes [6, 29, 16, 17] are extended to more complex 3D robotic manipulation tasks, and also achieve superior performance to other traditional architectures. However, applying diffusion models to 3D robotic manipulation faces the issue of insufficient decision-making frequency. 3D tasks involve complex spatial representations, which pose significant challenges for real-time robotic operation and control. Denoising processes are effective at generating high-fidelity outputs, but their iterative nature can be computationally expensive, creating a conflict between thorough exploration of the solution space and the need for fast decision-making in dynamic robotic tasks. To solve this issue, recent methods present to skip inference steps by hierarchical sampling [25, 15, 16], while it is hard to determine the hierarchy across different domains. In this paper, we harness the power of the consistency model [26] to improve the efficiency of diffusion models in 3D decision-making, utilizing point clouds as effective 3D representations.
II-B Consistency Models
@@
To accelerate the sampling speed of the current diffusion models (DMs) [30, 28, 31, 32], the concept of consistency models (CMs) [26] was first introduced in the image generation domain. The core idea behind the consistency models is to enforce the model to learn to map any point at any time to the origin of the probability flow ODE (PF-ODE) trajectory, i.e., the clean image. The consistency model facilitates efficient image generation, allowing trade-offs between extremely few inference steps and the generation quality. Recently, latent consistency models (LCMs) [27, 33] successfully distilled from the Stable Diffusion [32], achieving significant acceleration in the speed of text-conditioned image generation. Based on LCMs, [34, 35] introduce diverse conditional controls to manipulate the text-to-image latent consistency models. Moreover, the consistent model has been introduced in various research domains such as video [36, 37], 3D human motion [38], and audio [39], unlocking new capabilities for various real-time applications. While initial success has been achieved in incorporating relatively simple embodied tasks with the consistency model, as explored in [40, 41, 42], its application to complex robotic manipulation tasks with high-dimensional 3D visual conditions still remains largely unexplored. To fill this research gap, we propose ManiCM to achieve real-time 3D diffusion policy generation for the first time.
@@
III Approach
In this section, we first present the problem formulation and the preliminaries on consistency models (Section III-A). Then, we present an overview of our pipeline, and detail the design of each component in the manipulation consistency model (Section III-B). To enable our model to learn consistent generation from the teacher network, we propose the manipulation consistency distillation (Section III-C).
III-A Problem Formulation
The demand for robotic manipulation is of great significance in the pursuit of the general embodied agent. To train a manipulation agent, we focus on imitation learning that learns a policy from a limited set of expert demonstrations. The policy $\pi:\mathcal{O}\longmapsto\mathcal{A}$ maps the observation $o\in\mathcal{O}$ to the robot action $\boldsymbol{a}\in\mathcal{A}$. The observation contains the point cloud received from an eye-to-hand RGB-D camera and the robot proprioception data. ManiCM uses sparse point clouds as the 3D representation. Based on the 3D visual input $o^{(t)}$ and the robot poses, the agent is required to generate the optimal action for the robot arm and grippers $\boldsymbol{a}^{(t)}$, which respectively demonstrates the change in 3D space of the end-effector followed by a normalized torque that the gripper fingers should apply. While existing diffusion models produce high-quality robot actions, their iterative inference limits real-time 3D manipulation. Therefore, we propose a framework based on the consistency model to enhance decision efficiency.
@@ III-A Problem Formulation
As shown in Equation 1, the implication of the self-consistency property is that for any input pairs $(\mathbf{x}_{t},t)$ on the same PF-ODE trajectory, their outputs $\textbf{{{f}}}(\mathbf{x}_{t},t)$ remain consistent. All consistency models have to meet the boundary condition $\textbf{{{f}}}(\mathbf{x}_{\epsilon},\epsilon)=\mathbf{x}_{\epsilon}$, as it plays an important role in the success of consistency model training, according to [26]. To impose the self-consistency property on deep neural networks, we implement function $\textbf{{f}}(\cdot,\cdot)$ as:
@@
where $d(\cdot,\cdot)$ is a metric function chosen for measuring the distance between two samples. $\textbf{{{f}}}_{\bf\theta}(\cdot,\cdot)$ and $\textbf{{{f}}}_{{\bf\theta}^{-}}(\cdot,\cdot)$ are referred to as ‘online network’ and ‘target network’ according to [26]. When the optimization of the online network converges, the target network will eventually match the online network since $\bf{\theta}^{-}$ is a running average of $\bf{\theta}$.
III-B ManiCM: Manipulation Consistency Model
3D Conditional Generation. To incorporate the rich spatial information embedded in the 3D point cloud, we propose to inject the point cloud representation into the diffusion policy as conditions. Specifically, given the depth image received from a single-view RGB-D camera, we first obtain the point cloud by the extrinsic and intrinsic of the camera. We down-sample the point cloud via farthest point sampling, in order to focus on the interest region like the robot arm and targets. Then, the sparse point cloud is encoded by a carefully designed MLP point cloud encoder, resulting in a compact 3D representation. Subsequently, we condition the original diffusion model with the compact 3D representation, which can be expressed as:
@@ III-B ManiCM: Manipulation Consistency Model
where $\boldsymbol{\epsilon}_{\theta}$ is the predicted noise. $p$ and $q$ denote point cloud conditions and robot poses, respectively.
Table I: Comparisons on Runtime. We evaluate $100$ episodes on $31$ challenging tasks from Adroit and Metaworld across $3$ random seeds and report the time consumption per step (s) with standard deviation. The second results are underlined and the best results are bold. ‘’ denotes the reproduced version. The performance of our ManiCM in one-step inference surpasses all state-of-the-art models, providing ample evidence for the effectiveness of consistency distillation.
DP $10$ $162.6\scriptstyle{\pm 1.2}$
Image-based ManiCM ($1$-step) $1$ $\mathbf{14.9}$ $\mathbf{16.2}$ $\mathbf{16.1}$ $\mathbf{15.5}$ $\mathbf{16.3}$ $\mathbf{16.0\scriptstyle{\pm 0.1}}$
Voxel-based DP $10$ $183.6$ $205.3$ $192.2$ $198.8$ $211.5$ $202.1\scriptstyle{\pm 1.1}$
Voxel-based ManiCM ($1$-step) $1$ $\mathbf{18.1}$ $\mathbf{20.0}$ $\mathbf{20.6}$ $\mathbf{20.5}$ $\mathbf{20.8}$ $\mathbf{20.0\scriptstyle{\pm 0.2}}$
BC-GMM+3D $1$ $14.7$ $15.9$ $16.1$ $15.8$ $16.9$ $15.9\scriptstyle{\pm 0.7}$
BC-RNN+3D $1$ $14.9$ $16.8$ $16.2$ $16.4$ $16.4$ $16.6\scriptstyle{\pm 0.5}$
DP3 $10$ $159.6$ $177.4$ $186.1$ $183.2$ $181.8$ $177.6\scriptstyle{\pm 1.4}$
Simple DP3 $10$ $95.5$ $123.5$ $119.2$ $124.6$ $124.7$ $120.6\scriptstyle{\pm 0.9}$
DP3 ($4$-step) $4$ $52.6$ $66.5$ $76.3$ $58.6$ $62.1$ $64.9\scriptstyle{\pm 0.9}$
Simple DP3 ($4$-step) $4$ $27.3$ $34.4$ $33.6$ $35.6$ $34.2$ $33.6\scriptstyle{\pm 0.5}$
ManiCM ($4$-step) $4$ $56.3$ $66.8$ $68.2$ $67.9$ $69.4$ $66.2\scriptstyle{\pm 0.3}$
Table II: Comparisons on Success Rate. We evaluate $100$ episodes on $31$ challenging tasks from Adroit and Metaworld across $3$ random seeds and report the success rates (%) with standard deviation. The second results are underlined and the best results are bold. ‘’ denotes the reproduced version. The performance of our ManiCM in one-step inference surpasses all state-of-the-art models, providing ample evidence for the effectiveness of consistency distillation.
DP $10$ $31.6$ $77.1$ $18.0$ $3.0$ $\mathbf{22.5}$ $52.8\scriptstyle{\pm 3.6}$
Image-based ManiCM ($1$-step) $1$ $\mathbf{32.2}$ $\mathbf{77.5}$ $\mathbf{18.9}$ $\mathbf{4.3}$ $21.5$ $\mathbf{53.1\scriptstyle{\pm 3.9}}$
Voxel-based DP $10$ $35.7$ $80.2$ $24.5$ $\mathbf{12.4}$ $30.8$ $57.5\scriptstyle{\pm 3.1}$
Voxel-based ManiCM ($1$-step) $10$ $\mathbf{36.4}$ $\mathbf{80.5}$ $\mathbf{24.9}$ $11.5$ $\mathbf{31.5}$ $\mathbf{57.8\scriptstyle{\pm 3.8}}$
BC-GMM+3D $1$ $3.8$ $4.7$ $3.8$ $3.4$ $3.4$ $4.2\scriptstyle{\pm 0.4}$
BC-RNN+3D $1$ $4.1$ $4.5$ $3.8$ $3.2$ $3.0$ $4.0\scriptstyle{\pm 0.5}$
DP3 $10$ $76.1$ $91.7$ $46.3$ $44.8$ $62.3$ $77.5\scriptstyle{\pm 3.8}$
Simple DP3 $10$ $73.0$ $90.4$ $47.3$ $45.6$ $56.5$ $75.8\scriptstyle{\pm 3.3}$
DP3 ($4$-step) $4$ $75.3$ $86.1$ $39$ $44.0$ $61.3$ $72.4\scriptstyle{\pm 2.6}$
Simple DP3 ($4$-step) $4$ $73.0$ $84.3$ $37.3$ $43.3$ $62.0$ $70.9\scriptstyle{\pm 3.3}$
ManiCM ($1$-step) $1$ $74.9$ $92.9$ $47.7$ $44.2$ $\mathbf{65.7}$ $78.5\scriptstyle{\pm 4.2}$
@@
where $\textbf{{{F}}}_{\mathbf{\theta}}$ is a manipulation consistency function that contains the accumulated $\boldsymbol{\epsilon}_{\theta}$ in Equation 4 along the trajectory. In practice, $c_{\text{skip}}(t)=\sigma_{d}^{2}/(\beta^{2}t^{2}+\sigma_{d}^{2})$, $c_{\text{out}}(t)=\beta t/\sqrt{(\beta^{2}t^{2}+\sigma_{d}^{2})}$, where $\beta$ denotes a timestep scaling value while $\sigma_{d}$ is a balance value. The $\textbf{{{F}}}_{\mathbf{\theta}}(\cdot,\cdot)$ can be parametrized to predict the action sample or the noise epsilon. In mainstream diffusion generation works for image generation, the default prediction type is noise rather than sample. We speculate that while directly predicting samples may intuitively yield higher-quality generated outputs [43], predicting the distribution of images is much more challenging for denoising models in high-dimensional image space. However, the dimensionality of the robot action is quite small (approximately $28$) in manipulation tasks. In this scenario, directly learning the distribution of the samples is relatively easy and brings low variance and fast convergence, echoing [6]. More importantly, the large prediction variance caused by predicting noise will be amplified in consistency models. Specifically, the consistency property encourages the $\textbf{{{f}}}_{{\bf\theta}}(\boldsymbol{a}_{n},t_{n})$ to match $\hat{\boldsymbol{a}}_{0}=\textbf{{{f}}}_{{\bf\theta}}(\boldsymbol{a}_{0},t_{0})$, while the original diffusion model only requires $\hat{\boldsymbol{a}}_{n}$ to be close to $\hat{\boldsymbol{a}}_{n-1}$. Thus, the variance is propagated along the ODE trajectory in consistency modeling, which ultimately blocks the convergence of prediction noise in manipulation tasks (Figure 3). Therefore, we parameterize the manipulation self-consistency function $\textbf{{{F}}}_{\mathbf{\theta}}(\cdot,\cdot)$ to predict the action sample rather than noise.
III-C Manipulation Consistency Distillation